40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 18 21:37:59 2021, rana, Update, LSC, PRMI investigations: what IS the matrix?? PXL_20210319_045925024.jpg
    Reply  Fri Mar 19 11:18:25 2021, gautam, Update, LSC, PRMI investigations: what IS the matrix?? 
       Reply  Fri Mar 19 22:24:54 2021, gautam, Update, LSC, PRMI investigations: what IS the matrix?? IMG_9140.jpgIMG_9141.jpgREFL55.png
Message ID: 15942     Entry time: Thu Mar 18 21:37:59 2021     Reply to this: 15944
Author: rana 
Type: Update 
Category: LSC 
Subject: PRMI investigations: what IS the matrix?? 
  • Locked PRMI several tmes after Gautam setup. Easy w IFO CONFIG screenheart
  • tuned up alignment
  • Still POP22_I doesn't go above ~111, so not triggering the loops. Lowered triggers to 100 (POP22 units) and it locks fine now. smiley
  • Ran update on zita, and now it lost its mounts (and maybe its mind). Zita needs some love to recover the StripTool plots  sad
  • Put the $600 ebay TDS3052 near the LSC rack and tried to look at the RF power, but found lots of confusing information. Is there really a RF monitor in this demod board or was it disconnected by a crazy Koji cheeky ? I couldn't see any signal above a few mV.angry
  • Put a 20 dB coupler in line with the RF input and saw zip. Then I put the RF signal directly into the scope and saw that the 55 MHz signal is ~30 mVpp into 50 Ohms. I waited a few minutes with triggering to make sure I was getting the largest flashes. Why is the optical/RF signal so puny? surprise This is ~100x smaller than I think we want...its OK to saturate the RF stuff a little during lock acquisition as long as the loop can suppress it so that the RMS is < 3 dBm in the steady state.
Attachment 1: PXL_20210319_045925024.jpg  2.367 MB  Uploaded Thu Mar 18 23:13:31 2021  | Hide | Hide all
PXL_20210319_045925024.jpg
ELOG V3.1.3-